log5(x+2)-log5(10)= log5(100)
im confused on how to solve this perticular question.. if you could please help me id greatly apprectate it.. i was tryng to solve it different was...
ex.. log5(x+2)=100 + 2 = 102
102-10=92
10-2=8
my answers came up to be
108. 8. 998. 1000. 92 and i cant figure out what one is the actual answer.

Answer by lwsshak3(11628)

log5(x+2)-log5(10)= log5(100)
log5[(x+2)/10]= log5(100)
(x+2)/10=100
x+2=1000
x=998